simple series circuit inductor 1 henry capacitor of 10 power minus 6 Faradas and a resistor of 1000 ohms . the initial charge on a capacitor is zero if a 12 volt battery is connected to the circuit and the circuit is closed at t = 0 find the charge on the capacitor 1 second later and the steady state charge
1
Expert's answer
2020-05-19T08:59:40-0400

The voltage and the current relationships in circuit analysis

"i= \\frac{dq}{dt}\n\u200b"

voltage drop across the inductor

"V_L =L\\frac{di}{dt}=L\\frac{d^2q}{dt^2}\u200b"

voltage drop across the resistor

"V_R=Ri=R\frac{dq}{dt}"

voltage drop across the capacitor

"V _C=\frac{q}{C}"

Kirchhoff's Second Law

"L\frac{d^2q}{dt^2}+R\frac{dq}{dt}+\frac{q}{C}=E(t)"

Given that "L=1"H, "R=1000"ohm, "C=10^{-6}" F, "E(t)=12" V

"1\frac{d^2q}{dt^2}+1000\frac{dq}{dt}+\frac{1}{10^{-6}}q=12\\
\frac{d^2q}{dt^2}+1000\frac{dq}{dt}+1000000q=12\\"

The characteristic equation is

"r^2+1000r+1000000=0\\
D=(1000)^2-4\cdot1000000=-3000000\\
r_1=\frac{-1000+i1000\sqrt3}{2}=-500+500\sqrt3i\\
r_2=\frac{-1000-i1000\sqrt3}{2}=-500-500\sqrt3i\\
q_C=e^{-500t}(c_1\cos(500\sqrt3t)+c_2\sin(500\sqrt3t))"


Find the particular solution of the non-homogeneous differential equation using the method of the undetermined coefficients.

Assume that "Q(t)=A" is a solution of the non-homogeneous differential equation where "E(t)=12"

Differentiate the assumption with respect to "t"

"Q'=0, Q''=0"

Substitute in the original equation

"0+0+1000000A=12\\
A=1.2\cdot10^{-5}\\
Q(t)=1.2\cdot10^{-5}\\
q(t)= e^{-500t}(c_1\cos(500\sqrt3t)+c_2\sin(500\sqrt3t))+\\
+1.2\cdot10^{-5}"

The initial charge of the capacitor is zero

"q(0)= e^{-500\cdot 0}(c_1\cos(500\sqrt3\cdot 0)+\\
+c_2\sin(500\sqrt3\cdot 0))+1.2\cdot10^{-5}=0\\
c_1=-1.2\cdot10^{-5}"

The circuit is closed at "t=0: i(0)=q'(0)=0\\"

"q'(t)= -500e^{-500t}(-1.2\cdot10^{-5})(\cos(500\sqrt3t)-\\
-500e^{-500t}(c_2\sin(500\sqrt3t))+\\
+ 1.2\cdot10^{-5}(-500\sqrt3)e^{-500t}(\sin(500\sqrt3t)+\\
+c_2\cdot500\sqrt3e^{-500t}\cos(500\sqrt3t))\\
q'(0)=-500(-1.2\cdot10^{-5})+c_2\cdot500\sqrt3=0\\
c_2=-0.4\sqrt3\cdot10^{-5}\\
q(t)=e^{-500t}(-1.2\cdot10^{-5}\cos(500\sqrt3t)-\\
-0.4\sqrt3\cdot10^{-5}\sin(500\sqrt3t))+1.2\cdot10^{-5}"

Coulomb:

"q(1)=e^{-500\cdot(1)}(-1.2\cdot10^{-5}\cos(500\sqrt3\cdot(1))-\\
-0.4\sqrt3\cdot10^{-5}\sin(500\sqrt3\cdot(1)))+1.2\cdot10^{-5}\approx\\
\approx1.2\cdot10^{-5}(C)"

Find the steady state charge.

"t\to\infty:q_{p}(t)\to1.2\cdot10^{-5}C"